OpenStudy (anonymous):

Solve. 2x2 + x – 1 = 9

OpenStudy (anonymous):

let me post the answer choices!

OpenStudy (anonymous):

Do you me 2^2?

pooja195 (pooja195):

subtract the 9 so we have the equation = to 0 \[\huge 2x^2+x-10\] them multiply a*c  factor it out the factors are 5 and -4 they add to 1 and multiply to -20

pooja195 (pooja195):

(x+5)=0 (x-4)=0

pooja195 (pooja195):

then just solve the equations :) x+5=0 x-4=0

OpenStudy (mathmate):

Faster than the answer choices!

OpenStudy (anonymous):

the answer is 6

OpenStudy (anonymous):